OpenStudy (lilai3):

Please help! IT will be GREATLY appreciated! Will fan/award medal! i promise. you won't regret it (; The Waynesville auditorium seats 375 people. In a survey of 50 residents, 6 stated that they plan to attend the next town hall meeting. If the town has 4, 200 residents, how many would you expect to attend? Is the auditorium large enough?

Directrix (directrix):

Assuming that the 50 people surveyed are a random sample of the people in the community, then the following proportion would hold. 6/50 = x/4200 where x is the expected number to attend. 50*x = 6* 4200 x = ? And, will the auditorium be large enough?

OpenStudy (lilai3):

i got no the auditorium isn't large enough

Directrix (directrix):

Whatis the value of x? Without that, how do we know the auditorium is not large enough?

OpenStudy (lilai3):

x=504 the auditorium can only ho,ld 375 and 375<504

Directrix (directrix):

So, is the auditorium large enough if the expected number of people show up for the meeting?

OpenStudy (anonymous):

Key word here is 'expect' as in expected value. Very kind problem.

OpenStudy (lilai3):

no...???

Directrix (directrix):

Correct. You said it here: "the auditorium can only ho,ld 375 and 375<504" The thing is, you did not say how you knew which is why the expected value of x is needed.

OpenStudy (lilai3):

thanks for your great help!

Directrix (directrix):

It was fun. I don't regret it. :)

OpenStudy (lilai3):

lol